Recause of a recent wave ofJewelry store (d enes, a olyのaeases plice sveilance of jeuwelm stores. The inareosed sunsiltence costs chy on cetro 500,oo0 Per ucar, but as asotue amount of ewelny thatis stolen falls.Spechcalls without tta incresse in sorveillance ewen with a retel velve of S loo,ooo wald haue ben skolen Ts sien jeweiry wobld have been Renced bita euseiy thews r aucoo.ibhatis the ne sosial beneG+ Pom
Answer #1

Benefit to the society, B = Prevention of theft of stone robberies worth $ 1,000,000

However there is an incremental cost of surveillance, C = $ 500,000

Assumption:

  1. Thieves have no social recognition in the society. So society need not worry about their costs and benefits.

Net social benefit = B - C = 1,000,000 - 500,000 = $ 500,000